OpenStudy (anonymous):

Please help if i dont pass this test i could fail!! Phillip received 75 points on a project for school. He can make changes and receive two-tenths of the missing points back. He can make corrections as many times as he wants. Create the formula for the sum of this geometric series and explain your steps in solving for the maximum grade Phillip can receive. Identify this as converging or diverging.

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

@satellite73

OpenStudy (tkhunny):

I don't understand the question. Does he KNOW which questions are right or wrong? Does he fix them at the same rate as the original attempt? Too many questions for a rational answer.

OpenStudy (anonymous):

this is legit word for word the question and i dont get it at all

OpenStudy (tkhunny):

I agree with your assessment. 75 + (2/10)25 = 75 + 5 = 80 80 + (2/10)20 = 80 + 4 = 84 84 + (2/10)16 = 84 + 16/5 = 87 + 1/5

OpenStudy (anonymous):

would that be the answer?